Ratio and Proportion
3. An administrator of a popular web site is told that a new server can handle 41,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 105,000 hits per second; but every month, the peak demand increases by 2800 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.
Again, I am stumped on word problems. I had 24 questions and the last three are word problems that totally gets me confused because I don't know where to begin and what information I need to get started. Can someone please help me with this? It would be greatly appreciated!

Answer by stanbon(75887)

An administrator of a popular web site is told that a new server can handle 41,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 105,000 hits per second; but every month, the peak demand increases by 2800 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.
--------------
[# of servers/hits per sec] = [# of servers/hits per sec]
1/41000 = x/[105,000+24*2800]
x = [105,000+24*2800]/41000
-------
x = 172200/41000
x = 4.2
Rounding up they will need 5 servers to handle the peak demandin 24 mths.
-------------
How many servers do they have NOW?
105000/41000 = 2.56
So, they must have 3 servers NOW.
----------
Ans: they will need 2 new servers in 24 months.
=================================
Cheers,
Stan H.